How is ethane prepared from ethanol ????     Explain the reaction with the help of reaction
 

Grade:10

3 Answers

Arun
6 years ago
Ethanol can be converted to ethene in presence of concentrated sulphuric acid. It is called as dehydration reaction. 
CH​3-CH2-OH Concentrated H2SO4 CH2=CH2 + H2

b) Ethene can be converted to ethane by addition of hydrogen molecule. It is known as hydrogenation reaction. 
CH2=CH2 + H2 Nickel catalyst CH3-CH
Yash Chourasiya
3 years ago
Dear Student

1) Ethanol (C2H5OH) is dehydrated in the presence of excess conc. sulphuric acid to produce ethene (C2H4). It is called Dehydration reaction.
CH​3-CH2-OH →CH2=CH2+ H2O

2)Ethene upon hydrogenation in the presence of a catalyst like nickel or palladium yields ethane.It is known as hydrogenation reaction.
CH2=CH2+ H2 → CH3-CH3

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya
RISHIKA
3 years ago
HEY FRIEND
 Ethanol (C2H5OH) is dehydrated in the presence of excess conc. sulphuric acid to produce ethene (C2H4). It is called Dehydration reaction.
Ethene upon hydrogenation in the presence of a catalyst like nickel or palladium yields ethane.It is known as hydrogenation reaction.
THANKS